# BEND: Benchmarking DNA Language Models on Biologically Meaningful Tasks

### v1

@@ A.8 Extended results
CLL 0.87 0.81 0.80 0.86 0.82 0.84 0.83 0.85 0.88 0.86 0.83 0.86 0.84 0.86
Fibrobl 0.71 0.68 0.67 0.71 0.69 0.71 0.69 0.70 0.72 0.71 0.70 0.71 0.70 0.71
FibroP 0.78 0.69 0.65 0.75 0.70 0.74 0.72 0.74 0.77 0.75 0.72 0.75 0.72 0.76
Gliobla 0.86 0.75 0.71 0.84 0.77 0.82 0.80 0.82 0.86 0.83 0.80 0.83 0.81 0.85
GM12891 0.89 0.83 0.81 0.87 0.83 0.85 0.84 0.86 0.89 0.87 0.84 0.87 0.85 0.88
GM12892 0.88 0.84 0.83 0.87 0.85 0.86 0.85 0.86 0.89 0.87 0.86 0.87 0.86 0.88
GM18507 0.87 0.77 0.74 0.84 0.77 0.80 0.78 0.82 0.87 0.83 0.79 0.84 0.80 0.85
GM19238 0.86 0.79 0.77 0.84 0.80 0.82 0.81 0.83 0.87 0.84 0.81 0.84 0.82 0.85
GM19239 0.87 0.79 0.77 0.85 0.80 0.82 0.81 0.84 0.88 0.85 0.81 0.85 0.82 0.86
GM19240 0.81 0.74 0.73 0.80 0.76 0.77 0.76 0.78 0.82 0.79 0.77 0.80 0.77 0.80
H9ES 0.88 0.81 0.79 0.86 0.82 0.85 0.85 0.85 0.89 0.85 0.83 0.86 0.84 0.87
HeLa-S3_IFNa4h 0.85 0.72 0.70 0.82 0.76 0.81 0.79 0.81 0.85 0.81 0.79 0.82 0.79 0.83
Hepatocytes 0.72 0.73 0.72 0.75 0.74 0.75 0.74 0.74 0.76 0.76 0.75 0.76 0.74 0.75
HPDE6-E6E7 0.90 0.75 0.70 0.85 0.77 0.83 0.81 0.84 0.88 0.84 0.81 0.85 0.82 0.87
HSMM_emb 0.90 0.80 0.77 0.88 0.82 0.87 0.85 0.87 0.90 0.88 0.85 0.88 0.85 0.89
HTR8svn 0.91 0.76 0.72 0.86 0.78 0.84 0.82 0.85 0.89 0.85 0.82 0.86 0.83 0.88
Huh-7.5 0.81 0.76 0.75 0.81 0.78 0.80 0.79 0.80 0.83 0.81 0.79 0.81 0.79 0.82
Huh-7 0.84 0.77 0.75 0.83 0.78 0.81 0.80 0.81 0.86 0.82 0.80 0.82 0.80 0.84
iPS 0.91 0.87 0.87 0.90 0.88 0.90 0.89 0.90 0.91 0.90 0.88 0.90 0.89 0.91
Ishikawa_Estradiol 0.85 0.76 0.74 0.83 0.78 0.81 0.79 0.81 0.85 0.81 0.78 0.82 0.79 0.84
Ishikawa_4OHTAM 0.85 0.77 0.75 0.83 0.78 0.81 0.80 0.81 0.86 0.82 0.79 0.83 0.80 0.84
LNCaP_androgen 0.82 0.76 0.74 0.83 0.77 0.79 0.78 0.79 0.85 0.81 0.78 0.81 0.78 0.83
MCF-7_Hypoxia 0.83 0.75 0.74 0.81 0.76 0.79 0.78 0.80 0.85 0.80 0.77 0.80 0.78 0.81
Medullo 0.72 0.71 0.69 0.73 0.71 0.72 0.71 0.72 0.75 0.74 0.71 0.74 0.72 0.74
Melano 0.71 0.65 0.63 0.70 0.66 0.68 0.67 0.68 0.71 0.70 0.67 0.69 0.67 0.70
Myometr 0.84 0.74 0.68 0.82 0.75 0.80 0.78 0.80 0.84 0.81 0.77 0.81 0.79 0.83
Osteobl 0.72 0.69 0.68 0.72 0.70 0.72 0.71 0.71 0.73 0.72 0.71 0.72 0.71 0.73
PanIsletD 0.85 0.74 0.68 0.82 0.74 0.79 0.78 0.80 0.84 0.81 0.76 0.81 0.79 0.83
PanIslets 0.79 0.75 0.74 0.80 0.77 0.80 0.78 0.79 0.81 0.80 0.79 0.80 0.78 0.80
pHTE 0.81 0.73 0.70 0.79 0.74 0.78 0.76 0.78 0.81 0.78 0.76 0.78 0.77 0.80
ProgFib 0.85 0.76 0.71 0.83 0.76 0.80 0.79 0.81 0.85 0.82 0.78 0.82 0.80 0.84
RWPE1 0.90 0.74 0.68 0.85 0.76 0.83 0.81 0.83 0.88 0.84 0.80 0.84 0.82 0.87
Stellate 0.88 0.77 0.71 0.85 0.77 0.82 0.81 0.83 0.87 0.84 0.80 0.84 0.81 0.86
T-47D 0.81 0.75 0.73 0.79 0.75 0.78 0.77 0.78 0.81 0.79 0.77 0.79 0.76 0.80
CD4_Th0 0.79 0.76 0.75 0.80 0.77 0.79 0.78 0.79 0.80 0.80 0.78 0.79 0.78 0.80
Urothelia 0.90 0.79 0.76 0.87 0.81 0.85 0.84 0.86 0.89 0.87 0.83 0.86 0.84 0.88
Urothelia_UT189 0.85 0.76 0.73 0.82 0.78 0.81 0.80 0.81 0.88 0.82 0.79 0.81 0.80 0.84
AG04449 0.90 0.74 0.65 0.83 0.72 0.79 0.76 0.80 0.87 0.81 0.75 0.82 0.78 0.86
AG04450 0.89 0.74 0.66 0.84 0.74 0.80 0.78 0.81 0.87 0.82 0.77 0.83 0.79 0.86
AG09309 0.89 0.73 0.65 0.83 0.72 0.78 0.76 0.80 0.87 0.80 0.75 0.82 0.78 0.85
AG09319 0.89 0.75 0.67 0.84 0.74 0.80 0.78 0.81 0.87 0.82 0.77 0.83 0.79 0.85
AG10803 0.90 0.74 0.65 0.83 0.72 0.78 0.76 0.80 0.87 0.81 0.75 0.82 0.78 0.86
AoAF 0.89 0.74 0.66 0.83 0.73 0.79 0.77 0.80 0.87 0.81 0.76 0.82 0.78 0.85
BE2_C 0.80 0.73 0.69 0.80 0.72 0.76 0.75 0.77 0.83 0.78 0.73 0.79 0.75 0.81
BJ 0.89 0.75 0.66 0.83 0.73 0.79 0.77 0.80 0.87 0.81 0.76 0.82 0.78 0.85
Caco-2 0.91 0.91 0.90 0.92 0.91 0.91 0.91 0.91 0.93 0.92 0.90 0.92 0.91 0.92
CD20+ 0.87 0.78 0.73 0.84 0.76 0.80 0.79 0.83 0.88 0.84 0.78 0.84 0.80 0.85
CD34+ 0.87 0.75 0.70 0.83 0.73 0.78 0.77 0.82 0.87 0.81 0.76 0.83 0.78 0.84
CMK 0.81 0.74 0.69 0.81 0.72 0.76 0.74 0.79 0.85 0.78 0.72 0.81 0.76 0.81
GM06990 0.85 0.78 0.72 0.83 0.73 0.75 0.75 0.80 0.86 0.81 0.74 0.82 0.78 0.83
GM12864 0.85 0.74 0.67 0.80 0.69 0.73 0.72 0.78 0.85 0.79 0.71 0.80 0.76 0.82
GM12865 0.86 0.74 0.67 0.81 0.70 0.74 0.73 0.79 0.85 0.79 0.71 0.80 0.76 0.83
H7-hESC 0.78 0.72 0.69 0.78 0.73 0.75 0.75 0.76 0.83 0.77 0.71 0.79 0.75 0.79
HAc 0.88 0.74 0.66 0.83 0.74 0.79 0.78 0.81 0.87 0.81 0.76 0.83 0.78 0.86
HAEpiC 0.86 0.72 0.63 0.81 0.70 0.76 0.74 0.78 0.86 0.79 0.73 0.80 0.76 0.84
HA-h 0.88 0.74 0.67 0.83 0.74 0.79 0.78 0.81 0.86 0.81 0.77 0.83 0.79 0.85
HA-sp 0.84 0.73 0.68 0.81 0.73 0.78 0.76 0.79 0.83 0.80 0.76 0.80 0.77 0.82
HBMEC 0.89 0.73 0.64 0.83 0.72 0.80 0.78 0.82 0.87 0.82 0.77 0.83 0.79 0.86
HCF 0.89 0.74 0.66 0.83 0.73 0.79 0.77 0.80 0.87 0.81 0.76 0.83 0.78 0.86
HCFaa 0.89 0.72 0.64 0.83 0.72 0.80 0.77 0.81 0.87 0.81 0.76 0.82 0.78 0.85
HCM 0.89 0.73 0.65 0.82 0.72 0.78 0.76 0.80 0.87 0.80 0.75 0.82 0.77 0.85
HConF 0.89 0.74 0.67 0.84 0.74 0.81 0.78 0.81 0.87 0.82 0.78 0.83 0.79 0.86
HCPEpiC 0.88 0.71 0.64 0.81 0.71 0.77 0.75 0.79 0.85 0.79 0.74 0.80 0.76 0.84
HCT-116 0.89 0.73 0.68 0.85 0.75 0.84 0.81 0.85 0.88 0.84 0.80 0.85 0.81 0.87
HEEpiC 0.90 0.71 0.62 0.82 0.71 0.78 0.76 0.80 0.87 0.80 0.74 0.81 0.78 0.85
HFF 0.89 0.73 0.65 0.83 0.73 0.79 0.77 0.80 0.86 0.81 0.76 0.82 0.78 0.85
HFF-Myc 0.86 0.71 0.64 0.81 0.71 0.77 0.75 0.78 0.85 0.79 0.74 0.80 0.76 0.83
HGF 0.89 0.76 0.66 0.83 0.73 0.78 0.77 0.80 0.87 0.81 0.75 0.82 0.78 0.85
HIPEpiC 0.88 0.72 0.64 0.81 0.71 0.78 0.76 0.79 0.86 0.80 0.74 0.81 0.77 0.84
HL-60 0.81 0.70 0.64 0.77 0.67 0.69 0.68 0.73 0.83 0.76 0.66 0.75 0.72 0.80
HMF 0.90 0.74 0.64 0.84 0.73 0.81 0.78 0.82 0.88 0.83 0.77 0.83 0.80 0.87
HMVEC-dAd 0.89 0.76 0.70 0.85 0.76 0.81 0.79 0.83 0.88 0.83 0.79 0.85 0.80 0.87
HMVEC-dBl-Ad 0.89 0.74 0.66 0.84 0.72 0.78 0.76 0.82 0.88 0.81 0.75 0.83 0.78 0.86
HMVEC-dBl-Neo 0.88 0.73 0.66 0.82 0.72 0.78 0.76 0.81 0.86 0.80 0.75 0.82 0.77 0.85
HMVEC-dLy-Ad 0.88 0.76 0.68 0.83 0.74 0.79 0.78 0.81 0.87 0.81 0.77 0.83 0.79 0.86
HMVEC-dLy-Neo 0.89 0.75 0.67 0.84 0.73 0.79 0.77 0.82 0.87 0.81 0.76 0.83 0.79 0.86
HMVEC-dNeo 0.89 0.76 0.69 0.84 0.75 0.80 0.78 0.82 0.88 0.82 0.77 0.84 0.79 0.86
HMVEC-LBl 0.89 0.73 0.65 0.83 0.72 0.79 0.77 0.82 0.87 0.81 0.76 0.83 0.78 0.86
HMVEC-LLy 0.87 0.75 0.68 0.82 0.74 0.78 0.77 0.80 0.86 0.80 0.76 0.82 0.78 0.85
HNPCEpiC 0.89 0.72 0.64 0.83 0.72 0.79 0.77 0.81 0.87 0.81 0.76 0.82 0.78 0.85
HPAEC 0.88 0.75 0.68 0.83 0.74 0.80 0.78 0.82 0.87 0.82 0.77 0.84 0.79 0.86
HPAF 0.89 0.73 0.65 0.83 0.72 0.79 0.77 0.80 0.87 0.81 0.75 0.82 0.78 0.86
HPdLF 0.89 0.75 0.66 0.83 0.73 0.79 0.77 0.80 0.86 0.81 0.76 0.82 0.78 0.85
HPF 0.90 0.75 0.67 0.85 0.75 0.81 0.79 0.82 0.88 0.83 0.78 0.83 0.79 0.87
HRCEpiC 0.85 0.72 0.65 0.82 0.72 0.78 0.76 0.79 0.85 0.79 0.75 0.81 0.78 0.84
HRE 0.87 0.73 0.65 0.83 0.73 0.80 0.78 0.81 0.87 0.82 0.77 0.83 0.80 0.85
HRGEC 0.88 0.73 0.66 0.83 0.73 0.79 0.77 0.82 0.86 0.81 0.76 0.83 0.78 0.85
HRPEpiC 0.83 0.73 0.65 0.80 0.70 0.75 0.74 0.77 0.84 0.79 0.73 0.79 0.76 0.83
HVMF 0.86 0.74 0.65 0.82 0.71 0.77 0.75 0.78 0.85 0.80 0.74 0.80 0.76 0.83
Jurkat 0.82 0.72 0.65 0.80 0.67 0.71 0.70 0.78 0.84 0.76 0.68 0.78 0.74 0.82
Monocytes-CD14+ 0.86 0.74 0.67 0.82 0.71 0.75 0.74 0.80 0.88 0.81 0.72 0.81 0.77 0.84
NB4 0.87 0.74 0.68 0.83 0.72 0.77 0.75 0.80 0.88 0.81 0.74 0.82 0.77 0.85
NH-A 0.89 0.75 0.66 0.84 0.73 0.79 0.77 0.81 0.87 0.82 0.76 0.83 0.80 0.86
NHDF-Ad 0.87 0.74 0.64 0.81 0.70 0.76 0.75 0.78 0.85 0.79 0.73 0.80 0.76 0.84
NHDF-neo 0.87 0.76 0.65 0.82 0.71 0.77 0.76 0.79 0.86 0.80 0.74 0.81 0.77 0.84
NHLF 0.89 0.74 0.65 0.83 0.73 0.79 0.77 0.81 0.87 0.81 0.76 0.83 0.79 0.86
NT2-D1 0.82 0.74 0.71 0.81 0.76 0.79 0.78 0.80 0.85 0.80 0.76 0.82 0.78 0.82
PANC-1 0.86 0.71 0.65 0.82 0.73 0.81 0.79 0.82 0.85 0.81 0.78 0.82 0.78 0.84
PrEC 0.89 0.73 0.64 0.83 0.72 0.79 0.77 0.80 0.87 0.80 0.75 0.82 0.79 0.85
RPTEC 0.84 0.71 0.65 0.81 0.72 0.78 0.75 0.79 0.84 0.79 0.74 0.80 0.77 0.83
SAEC 0.90 0.71 0.62 0.82 0.71 0.79 0.77 0.80 0.87 0.80 0.75 0.81 0.78 0.85
SKMC 0.88 0.73 0.65 0.83 0.73 0.78 0.76 0.79 0.87 0.81 0.75 0.81 0.78 0.85
SK-N-MC 0.81 0.74 0.68 0.79 0.71 0.76 0.74 0.77 0.81 0.77 0.73 0.78 0.75 0.80
SK-N-SH_RA 0.87 0.85 0.81 0.89 0.83 0.85 0.85 0.86 0.90 0.87 0.83 0.88 0.85 0.89
Th2 0.86 0.79 0.73 0.84 0.76 0.78 0.78 0.83 0.87 0.82 0.77 0.84 0.81 0.86
WERI-Rb-1 0.75 0.75 0.65 0.81 0.70 0.70 0.72 0.77 0.86 0.79 0.67 0.79 0.75 0.82
WI-38 0.89 0.73 0.64 0.83 0.72 0.79 0.77 0.81 0.87 0.81 0.76 0.82 0.78 0.85
WI-38_4OHTAM 0.84 0.72 0.62 0.81 0.70 0.77 0.75 0.79 0.84 0.79 0.74 0.80 0.77 0.83
A549 0.84 0.71 0.67 0.81 0.74 0.79 0.78 0.80 0.84 0.80 0.77 0.80 0.78 0.82
GM12878 0.82 0.73 0.69 0.78 0.71 0.74 0.73 0.77 0.82 0.78 0.73 0.78 0.75 0.80
H1-hESC 0.86 0.82 0.80 0.85 0.82 0.84 0.84 0.84 0.87 0.84 0.82 0.85 0.84 0.86
HeLa-S3 0.82 0.70 0.66 0.79 0.71 0.76 0.75 0.77 0.82 0.78 0.74 0.78 0.76 0.80
HepG2 0.85 0.79 0.78 0.84 0.80 0.83 0.82 0.83 0.86 0.84 0.81 0.84 0.82 0.85
HMEC 0.80 0.71 0.68 0.77 0.72 0.76 0.74 0.76 0.79 0.76 0.73 0.77 0.75 0.78
HSMM 0.84 0.72 0.65 0.80 0.71 0.75 0.74 0.77 0.83 0.78 0.73 0.79 0.76 0.81
[…]

### v3

@@ A.8 Extended results
CLL 0.87 0.81 0.80 0.86 0.82 0.84 0.83 0.85 0.88 0.86 0.83 0.86 0.84 0.86 0.86
Fibrobl 0.71 0.68 0.67 0.71 0.69 0.71 0.69 0.70 0.72 0.71 0.70 0.71 0.70 0.71 0.71
FibroP 0.78 0.69 0.65 0.75 0.70 0.74 0.72 0.74 0.77 0.75 0.72 0.75 0.72 0.76 0.75
Gliobla 0.86 0.75 0.71 0.84 0.77 0.82 0.80 0.82 0.86 0.83 0.80 0.83 0.81 0.85 0.83
GM12891 0.89 0.83 0.81 0.87 0.83 0.85 0.84 0.86 0.89 0.87 0.84 0.87 0.85 0.88 0.87
GM12892 0.88 0.84 0.83 0.87 0.85 0.86 0.85 0.86 0.89 0.87 0.86 0.87 0.86 0.88 0.87
GM18507 0.87 0.77 0.74 0.84 0.77 0.80 0.78 0.82 0.87 0.83 0.79 0.84 0.80 0.85 0.84
GM19238 0.86 0.79 0.77 0.84 0.80 0.82 0.81 0.83 0.87 0.84 0.81 0.84 0.82 0.85 0.84
GM19239 0.87 0.79 0.77 0.85 0.80 0.82 0.81 0.84 0.88 0.85 0.81 0.85 0.82 0.86 0.85
GM19240 0.81 0.74 0.73 0.80 0.76 0.77 0.76 0.78 0.82 0.79 0.77 0.80 0.77 0.80 0.79
H9ES 0.88 0.81 0.79 0.86 0.82 0.85 0.85 0.85 0.89 0.85 0.83 0.86 0.84 0.87 0.86
HeLa-S3_IFNa4h 0.85 0.72 0.70 0.82 0.76 0.81 0.79 0.81 0.85 0.81 0.79 0.82 0.79 0.83 0.82
Hepatocytes 0.72 0.73 0.72 0.75 0.74 0.75 0.74 0.74 0.76 0.76 0.75 0.76 0.74 0.75 0.75
HPDE6-E6E7 0.90 0.75 0.70 0.85 0.77 0.83 0.81 0.84 0.88 0.84 0.81 0.85 0.82 0.87 0.85
HSMM_emb 0.90 0.80 0.77 0.88 0.82 0.87 0.85 0.87 0.90 0.88 0.85 0.88 0.85 0.89 0.88
HTR8svn 0.91 0.76 0.72 0.86 0.78 0.84 0.82 0.85 0.89 0.85 0.82 0.86 0.83 0.88 0.86
Huh-7.5 0.81 0.76 0.75 0.81 0.78 0.80 0.79 0.80 0.83 0.81 0.79 0.81 0.79 0.82 0.81
Huh-7 0.84 0.77 0.75 0.83 0.78 0.81 0.80 0.81 0.86 0.82 0.80 0.82 0.80 0.84 0.83
iPS 0.91 0.87 0.87 0.90 0.88 0.90 0.89 0.90 0.91 0.90 0.88 0.90 0.89 0.91 0.90
Ishikawa_Estradiol 0.85 0.76 0.74 0.83 0.78 0.81 0.79 0.81 0.85 0.81 0.78 0.82 0.79 0.84 0.82
Ishikawa_4OHTAM 0.85 0.77 0.75 0.83 0.78 0.81 0.80 0.81 0.86 0.82 0.79 0.83 0.80 0.84 0.83
LNCaP_androgen 0.82 0.76 0.74 0.83 0.77 0.79 0.78 0.79 0.85 0.81 0.78 0.81 0.78 0.83 0.82
MCF-7_Hypoxia 0.83 0.75 0.74 0.81 0.76 0.79 0.78 0.80 0.85 0.80 0.77 0.80 0.78 0.81 0.80
Medullo 0.72 0.71 0.69 0.73 0.71 0.72 0.71 0.72 0.75 0.74 0.71 0.74 0.72 0.74 0.73
Melano 0.71 0.65 0.63 0.70 0.66 0.68 0.67 0.68 0.71 0.70 0.67 0.69 0.67 0.70 0.69
Myometr 0.84 0.74 0.68 0.82 0.75 0.80 0.78 0.80 0.84 0.81 0.77 0.81 0.79 0.83 0.81
Osteobl 0.72 0.69 0.68 0.72 0.70 0.72 0.71 0.71 0.73 0.72 0.71 0.72 0.71 0.73 0.72
PanIsletD 0.85 0.74 0.68 0.82 0.74 0.79 0.78 0.80 0.84 0.81 0.76 0.81 0.79 0.83 0.82
PanIslets 0.79 0.75 0.74 0.80 0.77 0.80 0.78 0.79 0.81 0.80 0.79 0.80 0.78 0.80 0.80
pHTE 0.81 0.73 0.70 0.79 0.74 0.78 0.76 0.78 0.81 0.78 0.76 0.78 0.77 0.80 0.79
ProgFib 0.85 0.76 0.71 0.83 0.76 0.80 0.79 0.81 0.85 0.82 0.78 0.82 0.80 0.84 0.82
RWPE1 0.90 0.74 0.68 0.85 0.76 0.83 0.81 0.83 0.88 0.84 0.80 0.84 0.82 0.87 0.85
Stellate 0.88 0.77 0.71 0.85 0.77 0.82 0.81 0.83 0.87 0.84 0.80 0.84 0.81 0.86 0.84
T-47D 0.81 0.75 0.73 0.79 0.75 0.78 0.77 0.78 0.81 0.79 0.77 0.79 0.76 0.80 0.79
CD4_Th0 0.79 0.76 0.75 0.80 0.77 0.79 0.78 0.79 0.80 0.80 0.78 0.79 0.78 0.80 0.79
Urothelia 0.90 0.79 0.76 0.87 0.81 0.85 0.84 0.86 0.89 0.87 0.83 0.86 0.84 0.88 0.87
Urothelia_UT189 0.85 0.76 0.73 0.82 0.78 0.81 0.80 0.81 0.88 0.82 0.79 0.81 0.80 0.84 0.83
AG04449 0.90 0.74 0.65 0.83 0.72 0.79 0.76 0.80 0.87 0.81 0.75 0.82 0.78 0.86 0.83
AG04450 0.89 0.74 0.66 0.84 0.74 0.80 0.78 0.81 0.87 0.82 0.77 0.83 0.79 0.86 0.83
AG09309 0.89 0.73 0.65 0.83 0.72 0.78 0.76 0.80 0.87 0.80 0.75 0.82 0.78 0.85 0.82
AG09319 0.89 0.75 0.67 0.84 0.74 0.80 0.78 0.81 0.87 0.82 0.77 0.83 0.79 0.85 0.83
AG10803 0.90 0.74 0.65 0.83 0.72 0.78 0.76 0.80 0.87 0.81 0.75 0.82 0.78 0.86 0.83
AoAF 0.89 0.74 0.66 0.83 0.73 0.79 0.77 0.80 0.87 0.81 0.76 0.82 0.78 0.85 0.83
BE2_C 0.80 0.73 0.69 0.80 0.72 0.76 0.75 0.77 0.83 0.78 0.73 0.79 0.75 0.81 0.79
BJ 0.89 0.75 0.66 0.83 0.73 0.79 0.77 0.80 0.87 0.81 0.76 0.82 0.78 0.85 0.83
Caco-2 0.91 0.91 0.90 0.92 0.91 0.91 0.91 0.91 0.93 0.92 0.90 0.92 0.91 0.92 0.93
CD20+ 0.87 0.78 0.73 0.84 0.76 0.80 0.79 0.83 0.88 0.84 0.78 0.84 0.80 0.85 0.84
CD34+ 0.87 0.75 0.70 0.83 0.73 0.78 0.77 0.82 0.87 0.81 0.76 0.83 0.78 0.84 0.83
CMK 0.81 0.74 0.69 0.81 0.72 0.76 0.74 0.79 0.85 0.78 0.72 0.81 0.76 0.81 0.81
GM06990 0.85 0.78 0.72 0.83 0.73 0.75 0.75 0.80 0.86 0.81 0.74 0.82 0.78 0.83 0.82
GM12864 0.85 0.74 0.67 0.80 0.69 0.73 0.72 0.78 0.85 0.79 0.71 0.80 0.76 0.82 0.80
GM12865 0.86 0.74 0.67 0.81 0.70 0.74 0.73 0.79 0.85 0.79 0.71 0.80 0.76 0.83 0.81
H7-hESC 0.78 0.72 0.69 0.78 0.73 0.75 0.75 0.76 0.83 0.77 0.71 0.79 0.75 0.79 0.79
HAc 0.88 0.74 0.66 0.83 0.74 0.79 0.78 0.81 0.87 0.81 0.76 0.83 0.78 0.86 0.83
HAEpiC 0.86 0.72 0.63 0.81 0.70 0.76 0.74 0.78 0.86 0.79 0.73 0.80 0.76 0.84 0.81
HA-h 0.88 0.74 0.67 0.83 0.74 0.79 0.78 0.81 0.86 0.81 0.77 0.83 0.79 0.85 0.83
HA-sp 0.84 0.73 0.68 0.81 0.73 0.78 0.76 0.79 0.83 0.80 0.76 0.80 0.77 0.82 0.80
HBMEC 0.89 0.73 0.64 0.83 0.72 0.80 0.78 0.82 0.87 0.82 0.77 0.83 0.79 0.86 0.83
HCF 0.89 0.74 0.66 0.83 0.73 0.79 0.77 0.80 0.87 0.81 0.76 0.83 0.78 0.86 0.83
HCFaa 0.89 0.72 0.64 0.83 0.72 0.80 0.77 0.81 0.87 0.81 0.76 0.82 0.78 0.85 0.82
HCM 0.89 0.73 0.65 0.82 0.72 0.78 0.76 0.80 0.87 0.80 0.75 0.82 0.77 0.85 0.82
HConF 0.89 0.74 0.67 0.84 0.74 0.81 0.78 0.81 0.87 0.82 0.78 0.83 0.79 0.86 0.83
HCPEpiC 0.88 0.71 0.64 0.81 0.71 0.77 0.75 0.79 0.85 0.79 0.74 0.80 0.76 0.84 0.81
HCT-116 0.89 0.73 0.68 0.85 0.75 0.84 0.81 0.85 0.88 0.84 0.80 0.85 0.81 0.87 0.85
HEEpiC 0.90 0.71 0.62 0.82 0.71 0.78 0.76 0.80 0.87 0.80 0.74 0.81 0.78 0.85 0.81
HFF 0.89 0.73 0.65 0.83 0.73 0.79 0.77 0.80 0.86 0.81 0.76 0.82 0.78 0.85 0.82
HFF-Myc 0.86 0.71 0.64 0.81 0.71 0.77 0.75 0.78 0.85 0.79 0.74 0.80 0.76 0.83 0.81
HGF 0.89 0.76 0.66 0.83 0.73 0.78 0.77 0.80 0.87 0.81 0.75 0.82 0.78 0.85 0.83
HIPEpiC 0.88 0.72 0.64 0.81 0.71 0.78 0.76 0.79 0.86 0.80 0.74 0.81 0.77 0.84 0.81
HL-60 0.81 0.70 0.64 0.77 0.67 0.69 0.68 0.73 0.83 0.76 0.66 0.75 0.72 0.80 0.77
HMF 0.90 0.74 0.64 0.84 0.73 0.81 0.78 0.82 0.88 0.83 0.77 0.83 0.80 0.87 0.84
HMVEC-dAd 0.89 0.76 0.70 0.85 0.76 0.81 0.79 0.83 0.88 0.83 0.79 0.85 0.80 0.87 0.84
HMVEC-dBl-Ad 0.89 0.74 0.66 0.84 0.72 0.78 0.76 0.82 0.88 0.81 0.75 0.83 0.78 0.86 0.83
HMVEC-dBl-Neo 0.88 0.73 0.66 0.82 0.72 0.78 0.76 0.81 0.86 0.80 0.75 0.82 0.77 0.85 0.82
HMVEC-dLy-Ad 0.88 0.76 0.68 0.83 0.74 0.79 0.78 0.81 0.87 0.81 0.77 0.83 0.79 0.86 0.83
HMVEC-dLy-Neo 0.89 0.75 0.67 0.84 0.73 0.79 0.77 0.82 0.87 0.81 0.76 0.83 0.79 0.86 0.83
HMVEC-dNeo 0.89 0.76 0.69 0.84 0.75 0.80 0.78 0.82 0.88 0.82 0.77 0.84 0.79 0.86 0.84
HMVEC-LBl 0.89 0.73 0.65 0.83 0.72 0.79 0.77 0.82 0.87 0.81 0.76 0.83 0.78 0.86 0.83
HMVEC-LLy 0.87 0.75 0.68 0.82 0.74 0.78 0.77 0.80 0.86 0.80 0.76 0.82 0.78 0.85 0.82
HNPCEpiC 0.89 0.72 0.64 0.83 0.72 0.79 0.77 0.81 0.87 0.81 0.76 0.82 0.78 0.85 0.83
HPAEC 0.88 0.75 0.68 0.83 0.74 0.80 0.78 0.82 0.87 0.82 0.77 0.84 0.79 0.86 0.83
HPAF 0.89 0.73 0.65 0.83 0.72 0.79 0.77 0.80 0.87 0.81 0.75 0.82 0.78 0.86 0.83
HPdLF 0.89 0.75 0.66 0.83 0.73 0.79 0.77 0.80 0.86 0.81 0.76 0.82 0.78 0.85 0.83
HPF 0.90 0.75 0.67 0.85 0.75 0.81 0.79 0.82 0.88 0.83 0.78 0.83 0.79 0.87 0.84
HRCEpiC 0.85 0.72 0.65 0.82 0.72 0.78 0.76 0.79 0.85 0.79 0.75 0.81 0.78 0.84 0.81
HRE 0.87 0.73 0.65 0.83 0.73 0.80 0.78 0.81 0.87 0.82 0.77 0.83 0.80 0.85 0.83
HRGEC 0.88 0.73 0.66 0.83 0.73 0.79 0.77 0.82 0.86 0.81 0.76 0.83 0.78 0.85 0.83
HRPEpiC 0.83 0.73 0.65 0.80 0.70 0.75 0.74 0.77 0.84 0.79 0.73 0.79 0.76 0.83 0.80
HVMF 0.86 0.74 0.65 0.82 0.71 0.77 0.75 0.78 0.85 0.80 0.74 0.80 0.76 0.83 0.81
Jurkat 0.82 0.72 0.65 0.80 0.67 0.71 0.70 0.78 0.84 0.76 0.68 0.78 0.74 0.82 0.79
Monocytes-CD14+ 0.86 0.74 0.67 0.82 0.71 0.75 0.74 0.80 0.88 0.81 0.72 0.81 0.77 0.84 0.82
NB4 0.87 0.74 0.68 0.83 0.72 0.77 0.75 0.80 0.88 0.81 0.74 0.82 0.77 0.85 0.83
NH-A 0.89 0.75 0.66 0.84 0.73 0.79 0.77 0.81 0.87 0.82 0.76 0.83 0.80 0.86 0.84
NHDF-Ad 0.87 0.74 0.64 0.81 0.70 0.76 0.75 0.78 0.85 0.79 0.73 0.80 0.76 0.84 0.81
NHDF-neo 0.87 0.76 0.65 0.82 0.71 0.77 0.76 0.79 0.86 0.80 0.74 0.81 0.77 0.84 0.82
NHLF 0.89 0.74 0.65 0.83 0.73 0.79 0.77 0.81 0.87 0.81 0.76 0.83 0.79 0.86 0.83
NT2-D1 0.82 0.74 0.71 0.81 0.76 0.79 0.78 0.80 0.85 0.80 0.76 0.82 0.78 0.82 0.81
PANC-1 0.86 0.71 0.65 0.82 0.73 0.81 0.79 0.82 0.85 0.81 0.78 0.82 0.78 0.84 0.82
PrEC 0.89 0.73 0.64 0.83 0.72 0.79 0.77 0.80 0.87 0.80 0.75 0.82 0.79 0.85 0.82
RPTEC 0.84 0.71 0.65 0.81 0.72 0.78 0.75 0.79 0.84 0.79 0.74 0.80 0.77 0.83 0.80
SAEC 0.90 0.71 0.62 0.82 0.71 0.79 0.77 0.80 0.87 0.80 0.75 0.81 0.78 0.85 0.82
SKMC 0.88 0.73 0.65 0.83 0.73 0.78 0.76 0.79 0.87 0.81 0.75 0.81 0.78 0.85 0.82
SK-N-MC 0.81 0.74 0.68 0.79 0.71 0.76 0.74 0.77 0.81 0.77 0.73 0.78 0.75 0.80 0.78
SK-N-SH_RA 0.87 0.85 0.81 0.89 0.83 0.85 0.85 0.86 0.90 0.87 0.83 0.88 0.85 0.89 0.88
Th2 0.86 0.79 0.73 0.84 0.76 0.78 0.78 0.83 0.87 0.82 0.77 0.84 0.81 0.86 0.84
WERI-Rb-1 0.75 0.75 0.65 0.81 0.70 0.70 0.72 0.77 0.86 0.79 0.67 0.79 0.75 0.82 0.80
WI-38 0.89 0.73 0.64 0.83 0.72 0.79 0.77 0.81 0.87 0.81 0.76 0.82 0.78 0.85 0.83
WI-38_4OHTAM 0.84 0.72 0.62 0.81 0.70 0.77 0.75 0.79 0.84 0.79 0.74 0.80 0.77 0.83 0.80
A549 0.84 0.71 0.67 0.81 0.74 0.79 0.78 0.80 0.84 0.80 0.77 0.80 0.78 0.82 0.81
GM12878 0.82 0.73 0.69 0.78 0.71 0.74 0.73 0.77 0.82 0.78 0.73 0.78 0.75 0.80 0.78
H1-hESC 0.86 0.82 0.80 0.85 0.82 0.84 0.84 0.84 0.87 0.84 0.82 0.85 0.84 0.86 0.85
HeLa-S3 0.82 0.70 0.66 0.79 0.71 0.76 0.75 0.77 0.82 0.78 0.74 0.78 0.76 0.80 0.78
HepG2 0.85 0.79 0.78 0.84 0.80 0.83 0.82 0.83 0.86 0.84 0.81 0.84 0.82 0.85 0.84
HMEC 0.80 0.71 0.68 0.77 0.72 0.76 0.74 0.76 0.79 0.76 0.73 0.77 0.75 0.78 0.77
HSMM 0.84 0.72 0.65 0.80 0.71 0.75 0.74 0.77 0.83 0.78 0.73 0.79 0.76 0.81 0.80
[…]